OpenStudy (anonymous):

sec(39°) in terms of its cofunction

OpenStudy (fibonaccichick666):

can you give me an example of what a cofunction is?

OpenStudy (fibonaccichick666):

or the definition, whichever you prefer

OpenStudy (studygurl14):

http://www.mathwords.com/c/cofunction_identities.htm

OpenStudy (fibonaccichick666):

@aaleena, I need to know if you understand that part of the question. Can you please respond so we can move forward?

jimthompson5910 (jim_thompson5910):

if f(x) is a cofunction of g(x), then f(x) = g(90 - x) in degree mode f(x) = g(pi/2 - x) in radian mode

jimthompson5910 (jim_thompson5910):

example: COsine is a CO-function of sine

jimthompson5910 (jim_thompson5910):

which is why sin(x) = cos(90-x) and cos(x) = sin(90-x)

OpenStudy (fibonaccichick666):

@aaleena , if you scroll down (underneath this comment) there is an empty box, then click the little blue "post" button under on the right

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

yay! cool, so, now, do you understand the word cofunction?

OpenStudy (anonymous):

I know the answer to this question which is csc 51 degree But don't know how to show the work.

OpenStudy (fibonaccichick666):

ok, I can help you with that if you bear with me :)

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

so, do you know the cofunction for secant?

OpenStudy (anonymous):

cosecant

OpenStudy (anonymous):

cosecant which is csc right?

OpenStudy (fibonaccichick666):

yeppers

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

ok, so now, unfortunately you have to memorize the identity here but if you look at jim's post, you will see the basic idea. Let me know if you understand his explanation for degrees

OpenStudy (anonymous):

okay but where is Jim's post?

OpenStudy (fibonaccichick666):

scroll up a little and you will se a guy with the username @jim_thompson5910

OpenStudy (anonymous):

never mind ... I just found Jim's post

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

okey dokey, so now, let me know if you understand the first part with "degree mode"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yeah I can understand it

OpenStudy (fibonaccichick666):

ok, so how we get between uses that, we have sec(39) In our cofunction, we can use csc(90-x) So, our x=39, can you substitute that into the cosecant for me please?

OpenStudy (anonymous):

csc(90-39) = 51

OpenStudy (fibonaccichick666):

yep, so that is how you get csc(51)

OpenStudy (anonymous):

okay. I have another question how can I solve this sin 1/2

OpenStudy (fibonaccichick666):

ok, make a new post. Close this question. Give someone a medal, then write a new question and tag me using the @ symbol and my username

OpenStudy (anonymous):

ok